Consider the following results of a simple regression model of dollar price of unleaded gas (dependent variable) and dollar price of crude oil (independent variable):

Coefficient

t-statistics

Intercept

0.56

28.27

Crude Oil

0.0457

73.34

R-square = 0.87

  1. What will be expected change in the price of unleaded gas if the crude oil price is expected to fall by 7%. The current price of crude oil and unleaded gas are $74 and $3.10 respectively?
  1. (7.89%)
  2. (6.85%)
  3. (7.25%)
  4. (7.64%)
  1. What will be forecasted price of unleaded gas if the crude oil price is expected to fall by 3%. The current price of crude oil and unleaded gas are $74 and $3.10 respectively?
  1. $3
  2. $2.93
  3. $2.86
  4. $3.07
  1. What is the interpretation of coefficient for crude oil?
  1. Every $1 increase in crude price will cause unleaded gas price to increase by 4.57 cents.
  2. Every $1 increase in crude price will cause unleaded gas price to increase by 4.57 %.
  3. Every 1% increase in crude price will cause unleaded gas price to increase by 4.57%.
  4. Every 1% increase in crude price will cause unleaded gas price to increase by 4.57 cents.
  1. What will be forecasted price of unleaded gas when the crude oil price is $80?
  1. $3.99
  2. $4.44
  3. $4.22
  4. $4.67
Answer #1

a.

The regression equation is,

Unleaded gas = 0.56 + 0.0457 crude oil price

If the crude oil price is expected to fall by 7%, then,

New Unleaded gas = 0.56 + 0.0457 crude oil price * (1 - 0.07)

Change in Price = New Unleaded gas - Unleaded gas

= (0.56 + 0.0457 crude oil price) - (0.56 + 0.0457 crude oil price * (1 - 0.07))

= 0.0457 crude oil price * 0.07

= 0.0457 * 74 * 0.07

= 0.236726

Expected percent change = (Change / Unleaded gas price) * 100

= (0.236726 / 3.10) * 100

= 7.636323 %

d. (7.64%)

b.

The regression equation is,

Unleaded gas = 0.56 + 0.0457 crude oil price

If the crude oil price is expected to fall by 3%, then,

  New Unleaded gas = 0.56 + 0.0457 crude oil price * (1 - 0.03)

Change in Price = New Unleaded gas - Unleaded gas

= (0.56 + 0.0457 crude oil price) - (0.56 + 0.0457 crude oil price * (1 - 0.03))

= 0.0457 crude oil price * 0.03

= 0.0457 * 74 * 0.03

= 0.101454

Forecasted price of unleaded gas = 3.10 - 0.101454

= 3

a. $3

c.

Slope of the regression line = 0.0457

Every $1 increase in crude price will cause unleaded gas price to increase by $0.0457 = 4.57 cents

Thus,

a. Every $1 increase in crude price will cause unleaded gas price to increase by 4.57 cents.

d.

The regression equation is,

Unleaded gas = 0.56 + 0.0457 crude oil price

For crude oil price = $80

Unleaded gas = 0.56 + 0.0457 * 80 = 4.216

c. $4.22